# A Survey of Foundation Models for IoT: Taxonomy and Criteria-Based Analysis

### v1

@@ 3. Criterion I: Efficiency
Efficient Transformer Architecture. The Transformer architecture (Vaswani et al., 2017) has been widely adopted in the IoT domain to build foundation models (e.g., IoT-LM (Mo et al., 2024) and LIMU-BERT (Xu et al., 2021)). However, Transformers are known for their high computational complexity, driven by three key factors: (1) the self-attention mechanism, which scales quadratically with input length; (2) a deep multilayer architecture with unshared trainable parameters across different layers; and (3) fully-connected feedforward layers, which contain the majority of the model’s parameters.
To address the computational challenges of self-attention, sparse attention mechanisms (Tay et al., 2022; Child et al., 2019; Ho et al., 2019; Chen et al., 2023; Wei et al., 2024b) can be employed. Rather than allowing each token to attend to all others, sparse attention restricts attention to a limited set of positions, significantly reducing computational complexity. Additionally, attention mechanisms can be optimized for hardware-specific constraints, such as GPU memory and I/O throughput (Dao et al., 2022; Dao, 2023). To mitigate the overhead of multi-layer architectures, LIMU-BERT (Xu et al., 2021) introduces cross-layer parameter sharing, where only the first encoder layer is trained and its parameters are reused across all subsequent layers. This approach drastically reduces the total number of parameters, improving efficiency and enabling deployment on resource-constrained devices. Finally, to reduce the computational cost of fully connected layers, Mixture-of-Experts (MoE) architectures (Cai et al., 2025) can be used. MoE layers employ sparse gating to activate only a small subset of expert sub-networks per token, rather than processing each token through all experts. This design enables models to scale to billions of parameters while maintaining efficient training and inference. Recent work such as LiteMoE (Zhuang et al., 2024) further enhances the efficiency of MoE-based architectures for the on-device deployment.
@@ 4. Criterion II: Context Awareness
Context awareness refers to a system’s ability to dynamically adapt its behavior based on situational factors relevant to a specific task (Perera et al., 2013; Dey et al., 2000). In the context of applying FMs to IoT tasks, contextual factors typically fall into two key categories: (1) environmental context, which includes physical conditions (e.g., time, location, room brightness, available IoT devices) (Guo et al., 2024b; Chen et al., 2024b; Cui et al., 2024a) and virtual settings (e.g., accessible software applications) (Liu et al., 2024b); and (2) user-specific context, which accounts for individual preferences, schedules, backgrounds, and personalities.
@@
Context awareness is critical for effectively applying FMs to IoT tasks. First, it enables FMs to accurately interpret raw sensor data, which typically consists of unstructured numerical values. Since these values reflect real-world physical states, understanding their meaning requires contextual metadata (e.g., sampling rate, sensor placement, and measurement units). Without this information, FMs struggle to infer the true significance of the data, impairing reasoning and decision-making (An et al., 2024). Second, since FMs are pre-trained on broad datasets to capture general patterns, they often overlook context-specific cues essential for task performance in real-world environments (Yuan et al., 2023; Prasad et al., 2023). For instance, LLMs may generate overly generic outputs when they lack grounding in environmental context (e.g., availability of physical objects or tools), which limits their ability to produce actionable, situation-appropriate responses. Third, context-aware foundation models can dynamically adapt to diverse IoT scenarios without requiring separate models for each specific case. This adaptability is especially valuable in IoT environments, where tasks are inherently context-dependent and conditions frequently change. By enabling a single model to generalize across varying situations, context awareness enhances system flexibility and automation, while also reducing the burden of developing and maintaining multiple specialized models. Finally, incorporating user-specific context (e.g., location, activity, or preferences) allows FM-based IoT systems to personalize services and interactions. This not only improves usability but also enhances user engagement and satisfaction (Yang et al., 2025).
@@
The collected papers propose various methods to enable context awareness in foundation models for IoT applications, which we summarize below.
@@
Data Enrichment. Data enrichment involves augmenting formatted and encoded sensor inputs with additional contextual information to help foundation models more accurately interpret the underlying physical meaning behind the data. Critical background details (e.g., units of measurement, sampling rate, and device placement) are essential, as identical sensor readings can have vastly different meanings depending on the context (An et al., 2024; Ji et al., 2024). For instance, sequences of the same length may reflect different durations depending on sampling rates, and identical values could represent entirely different phenomena (e.g., movement speed vs. body temperature) based on the measurement unit. Without this contextual information, foundation models may misinterpret the data and make suboptimal decisions. Therefore, such metadata should be included alongside raw sensor inputs when using prompt-based, agent-based, or training-based approaches.
@@
Fine-Tuning. One of the most effective ways to achieve context awareness in foundation models, originally trained on general-purpose datasets, is to fine-tune them using data specific to the target environment, user, and task. This process adjusts the model’s parameters to capture context-specific patterns while preserving the general knowledge gained during pretraining. Fine-tuning significantly improves the model’s ability to interpret and respond to contextual cues. For instance, Guo et al. (2024b) fine-tuned an LLM for an augmented reality (AR) system using user interaction and feedback data, enabling the model to generate virtual scenes more closely aligned with individual user preferences.
@@
Contextual Inputs. While fine-tuning can significantly enhance the context awareness of foundation models, it often requires substantial computational resources and large volumes of training data. A more lightweight alternative when applying prompt-based or agent-based methods is to incorporate environmental and personal context directly into the model inputs as textual prompts. For example, Nepal et al. (2024) personalize a journaling system by embedding user context, such as current mood, stress level, and summaries of daily behavior (e.g., screen time, walking duration), into the prompt before passing it to the LLM for reasoning. Similarly, LLMind (Cui et al., 2024a) incorporates user profiles (e.g., background and experience) into the system prompt and leverages the LLM’s role-playing capabilities (Chen et al., 2024c) to generate tailored responses.
@@
However, implementing context-aware personalization may raise significant privacy concerns. Developers must be careful of these issues and avoid intrusive practices (e.g., collecting personal information from users’ social media without explicit consent) (Yang et al., 2025).
@@
Contextual Memory. Contextual information can also be stored in external memory to enhance the context awareness when applying agent-based methods. This external repository should be updated in real time or at regular intervals to remain adaptive to contextual changes.
@@
Several studies have proposed storing environmental or personal data in such memory structures. For example, Yang et al. (2025) uses an LLM to extract persona-related information (e.g., personal interests, experiences, and background) from conversations and stores it in a persona database at the end of each interaction. For new conversational partners, the extracted persona are directly registered into the memory. For known users or previously encountered partners, they use the LLM to check for existing persona in the database. If no conflicting or redundant information is found, the extracted new persona are added. If the persona are semantically similar to existing entries, they are merged. If contradictions are detected, the old information in the database are replaced by the new ones. The stored personal information can then be integrated into the LLM’s reasoning, either by accessing it through external memory during inference or by incorporating relevant context into prompts, enhancing the model’s adaptability to individual users.
@@
Active Sensing. In addition to passively receiving environmental data (e.g., sensor data) or user information (e.g., through user interactions), agent-based methods can also actively query physical IoT devices (e.g., sensors, robots) to gather additional information when the current context is insufficient for high-level reasoning or decision-making. This is referred to as the active sensing capability. For example, in the LLMind (Cui et al., 2024a) paper, an agent-based check-in and security system can prompt a robot to approach and identify a person when the system cannot recognize them from low-resolution images captured by a ceiling-mounted camera or due to occlusion.
@@
Iterative Self-Refinement. For methods based on LLMs, due to their non-deterministic nature, incorporating contextual information into prompts does not always guarantee fully executable responses or complete satisfaction of personalization requirements. To address this, iterative self-refinement mechanism are proposed to transforms LLMs into a closed-loop system (Liu et al., 2024b; Cui et al., 2024a; Guo et al., 2024b). Specifically, the LLM’s output (e.g., a plan generated by an LLM-agent) receives feedback from the environment or the user if the response is not executable, possibly due to overlooked environmental or user-specific factors. The LLM is then reprompted with information about the execution error or user feedback, generating a revised response. This process repeats until the generated response fully meets the environmental and personalization requirements.
@@ 5. Criterion III: Safety
Safety in IoT systems refers to preventing the system and its components from causing physical harm or posing threats, and to protecting the surrounding environment from such risks (Atlam and Wills, 2020). It is worth noting that in other related surveys (e.g., Ma et al. (2025)), safety may also encompass security, which refers to protecting models from external threats such as adversarial, jailbreak attacks. In this survey, we consider them differently as Shi et al. (2024) and address security separately in Section 6.
@@
All studies reviewed in this survey focus on enhancing the safety of LLMs; therefore, we summarize the commonly used techniques for improving LLM safety in the context of IoT integration.
@@
Safety Requirement Incorporation. Embedding safety requirements into LLM inputs, typically via system prompts, is a key strategy for promoting safety awareness (Xu et al., 2024b). This approach requires developers and domain experts to proactively identify potential risks and define appropriate constraints and mitigation strategies prior to deployment. In agent-based systems, where external information can be retrieved dynamically from online sources or databases, environment- and task-specific safety requirements can also be automatically incorporated into the prompt to ensure context-aware and adaptive safety compliance.
@@
User Request Verification. When presented with a task from user requests, an LLM should reason over both the request and the current environmental context to assess whether the task is safe and compliant with predefined safety constraints. Context-aware reasoning is essential, as high-level instructions may not be safe to execute across all situations (Cui et al., 2024b; Xu et al., 2024b). For example, the model should recognize and reject unsafe user requests, such as driving at 200 miles per hour (322 km/h) on a snowy road.
@@
Post-Response Verification. After the LLM generates the response, it is crucial to verify that the output adheres to both general and task-specific safety constraints, either manually or automatically. Manual verification, however, can be time-consuming, especially when domain expertise and complex reasoning are required. Therefore, automated verification methods are essential. For example, Xu et al. (2024b) utilizes the Z3 Python API to check safety constraints formulated as first-order logic (FOL) against LLM-generated plans, which are expressed in formal languages such as Linear Temporal Logic (LTL) (Xu et al., 2024b; Yang et al., 2024b). Alternatively, responses can be validated through simulations to ensure safe and correct execution before deployment in real-world environments.
@@
Interpretable Decision-Making. When the LLM makes the decisions or operates the physical devices, especially in the high-stake scenarios (e.g., healthcare, autonomous driving), interpretability of the final decision is very important. This means being able to provide not only the final decision, but also the reasoning procedure explicitly (e.g., reasoning chain in Chain-of-Thought). This can help humans identify if the generated decision is correct and avoid catastrophes if the LLM intermediate reasoning steps violates the safety rules (Cui et al., 2024b).
@@
Iterative Self-Refinement. If an LLM output violates safety constraints, the model should be reprompted and required to regenerate the output iteratively, guided by feedback from either human reviewers or automated verification systems, until the generated response satisfies all predefined safety requirements (Xu et al., 2024b; Yang et al., 2024b).

### v2

@@ 4. Criterion I: Efficiency
Efficient Transformer Architecture. The Transformer architecture (Vaswani et al., 2017) has been widely adopted in the IoT domain to build foundation models (e.g., IoT-LM (Mo et al., 2024) and LIMU-BERT (Xu et al., 2021)). However, Transformers are known for their high computational complexity, driven by three key factors: (1) the self-attention mechanism, which scales quadratically with input length; (2) a deep multilayer architecture with unshared trainable parameters across different layers; and (3) fully-connected feedforward networks (FFNs), which contain the majority of the model’s parameters ($\geq 66\%$, typically) (Vaswani et al., 2017).
To address the computational challenges of self-attention, sparse attention mechanisms (Tay et al., 2022; Child et al., 2019; Ho et al., 2019; Chen et al., 2023; Wei et al., 2024b) can be employed. Rather than allowing each token to attend to all others, sparse attention restricts attention to a limited set of positions, significantly reducing computational complexity. However, sparse attention has not been extensively validated at scale in large foundation models, leaving an open question for future research. Furthermore, in question answering systems (e.g., IoT-LM (Mo et al., 2024) and Sensor2Text (Chen et al., 2024a)) that use autoregressive Multi-Head Attention (MHA) for text generation, the key-value (KV) cache is a widely adopted technique to avoid redundant attention computations during inference. To further reduce memory usage and computational overhead, several attention mechanisms have been proposed to share or compress key-value pairs, such as Multi-Query Attention (MQA) (Shazeer, 2019), Grouped-Query Attention (GQA) (Ainslie et al., 2023), and Multi-Head Latent Attention (MLA) (Liu et al., 2024a). Additionally, attention mechanisms can be optimized for hardware-specific constraints, such as GPU memory and I/O throughput (Dao et al., 2022; Dao, 2023).
To mitigate the overhead of multi-layer architectures, LIMU-BERT (Xu et al., 2021) introduces cross-layer parameter sharing, where only the first encoder layer is trained and its parameters are reused across all subsequent layers. This approach drastically reduces the total number of parameters, improving efficiency and enabling deployment on resource-constrained devices. Like sparse attention, cross-layer parameter sharing has yet to be extensively validated at scale in large foundation models, presenting an open question for future research.
Finally, to reduce the computational cost of FFN layers, Mixture-of-Experts (MoE) architectures (Cai et al., 2025) can be used. MoE layers employ sparse gating to activate only a small subset of expert sub-networks per token, rather than processing each token through all experts. This design enables models to scale to billions of parameters while maintaining efficient training and inference. Recent work such as LiteMoE (Zhuang et al., 2024) further enhances the efficiency of MoE-based architectures for the on-device deployment. However, MoE structures are typically not applied to other densely parameterized components of Transformer architectures, such as the projection matrices in attention layers.
@@ 5. Criterion II: Context Awareness
Context awareness refers to a system’s ability to dynamically adapt its behavior based on situational factors relevant to a specific task (Perera et al., 2013; Dey et al., 2000). In the context of applying FMs to IoT tasks, contextual factors typically fall into two key categories: (1) environmental context, which includes physical conditions (e.g., time, location, room brightness, available IoT devices) (Guo et al., 2024b; Chen et al., 2024b; Cui et al., 2024a) and virtual settings (e.g., accessible software applications) (Liu et al., 2024b); and (2) user-specific context, which accounts for individual preferences, schedules, backgrounds, and personalities (Yang et al., 2025a, b).
@@
Context awareness is critical for effectively applying FMs to IoT tasks. First, it enables FMs to accurately interpret raw sensor data, which typically consists of unstructured numerical values. Since these values reflect real-world physical states, understanding their meaning requires contextual metadata (e.g., sampling rate, sensor placement, and measurement units). Without this information, FMs struggle to infer the true significance of the data, impairing reasoning and decision-making (An et al., 2024). Second, since FMs are pre-trained on broad datasets to capture general patterns, they often overlook context-specific cues essential for task performance in real-world environments (Yuan et al., 2023; Prasad et al., 2023). For instance, LLMs may generate overly generic outputs when they lack grounding in environmental context (e.g., availability of physical objects or tools), which limits their ability to produce actionable, situation-appropriate responses. Third, context-aware foundation models can dynamically adapt to diverse IoT scenarios without requiring separate models for each specific case. This adaptability is especially valuable in IoT environments, where tasks are inherently context-dependent and conditions frequently change. By enabling a single model to generalize across varying situations, context awareness enhances system flexibility and automation, while also reducing the burden of developing and maintaining multiple specialized models. Finally, incorporating user-specific context (e.g., location, activity, or preferences) allows FM-based IoT systems to personalize services and interactions. This not only improves usability but also enhances user engagement and satisfaction (Yang et al., 2025a).
@@
The collected papers propose various methods to enable context awareness in foundation models for IoT applications, which we summarize below. Table 4 also provides an overview of these methods, including their descriptions and the paradigms (prompt-based, agent-based, training-based) to which they are mostly commonly applied.
Table 4. Summary of Context-Awareness-Enhancing Methods, with Descriptions and Typically Applicable Paradigms (Prompt-Based, Agent-Based, Training-Based).
Typically Applicable Paradigm
Method Description Prompt Agent Training
Data Enrichment
Augmenting formatted and encoded sensor inputs with additional
contextual information to help foundation models more accurately
interpret the underlying physical meaning behind the data.
Fine-Tuning
Fine-tuning pre-trained foundation models using data specific
to the target environment, user, and task.
Contextual Inputs
Incorporating environmental and personal context directly
into the model inputs as textual prompts.
Contextual Memory
Using long-term external memory to store kwledge
and information about the agent’s current context.
Active Sensing
Actively querying physical IoT devices (e.g., sensors, robots) to
gather additional information.
Iterative Self-Refinement Refining outputs based on feedback from environments or users.
@@
Data Enrichment. Data enrichment involves augmenting formatted and encoded sensor inputs with additional contextual information to help foundation models more accurately interpret the underlying physical meaning behind the data. Critical background details (e.g., units of measurement, sampling rate, and device placement) are essential, as identical sensor readings can have vastly different meanings depending on the context (An et al., 2024; Ji et al., 2024; Li et al., 2024b; Ouyang and Srivastava, 2024; Xu et al., 2024a). For instance, sequences of the same length may reflect different durations depending on sampling rates, and identical values could represent entirely different phenomena (e.g., movement speed vs. body temperature) based on the measurement unit. Without this contextual information, foundation models may misinterpret the data and make suboptimal decisions. Therefore, such metadata should be included alongside raw sensor inputs when using prompt-based, agent-based, or training-based approaches.
@@
Fine-Tuning. One of the most effective ways to achieve context awareness in foundation models, originally trained on general-purpose datasets, is to fine-tune them using data specific to the target environment, user, and task. This process adjusts the model’s parameters to capture context-specific patterns while preserving the general knowledge gained during pretraining (Guo et al., 2024b; Tian et al., 2025). Fine-tuning significantly improves the model’s ability to interpret and respond to contextual cues. For instance, Guo et al. (2024b) fine-tuned an LLM for an augmented reality (AR) system using user interaction and feedback data, enabling the model to generate virtual scenes more closely aligned with individual user preferences.
@@
Contextual Inputs. While fine-tuning can significantly enhance the context awareness of foundation models, it often requires substantial computational resources and large volumes of training data. A more lightweight alternative when applying prompt-based or agent-based methods is to incorporate environmental and personal context directly into the model inputs as textual prompts (Nepal et al., 2024; Tian et al., 2025; Cui et al., 2024a; Huang et al., 2022; Song et al., 2023b; Singh et al., 2022). For example, Nepal et al. (2024) personalize a journaling system by embedding user context, such as current mood, stress level, and summaries of daily behavior (e.g., screen time, walking duration), into the prompt before passing it to the LLM for reasoning. Similarly, LLMind (Cui et al., 2024a) incorporates user profiles (e.g., background and experience) into the system prompt and leverages the LLM’s role-playing capabilities (Chen et al., 2024c) to generate tailored responses.
@@
However, implementing context-aware personalization may raise significant privacy concerns. Developers must be careful of these issues and avoid intrusive practices (e.g., collecting personal information from users’ social media without explicit consent) (Yang et al., 2025a).
@@
Contextual Memory. Contextual information can also be stored in external memory (called contextual repository (Cui et al., 2024a) or semantic memory (Kumar, 2021; Nuxoll and Laird, 2007)) to enhance the context awareness when applying agent-based methods. This external repository should be updated in real time or at regular intervals to remain adaptive to contextual changes.
@@
Several studies (Rivkin et al., 2024; Cui et al., 2024a; Saleh et al., 2025) have proposed storing environmental or personal data in such memory structures. For example, Yang et al. (2025a) uses an LLM to extract persona-related information (e.g., personal interests, experiences, and background) from conversations and stores it in a persona database at the end of each interaction. For new conversational partners, the extracted persona are directly registered into the memory. For known users or previously encountered partners, they use the LLM to check for existing persona in the database. If no conflicting or redundant information is found, the extracted new persona are added. If the persona are semantically similar to existing entries, they are merged. If contradictions are detected, the old information in the database are replaced by the new ones. The stored personal information can then be integrated into the LLM’s reasoning, either by accessing it through external memory during inference or by incorporating relevant context into prompts, enhancing the model’s adaptability to individual users.
@@
Active Sensing. In addition to passively receiving environmental data (e.g., sensor data) or user information (e.g., through user interactions), agent-based methods can also actively query physical IoT devices (e.g., sensors, robots) to gather additional information when the current context is insufficient for high-level reasoning or decision-making (Cui et al., 2024a; Yang et al., 2025b). This is referred to as the active sensing capability. For example, in the LLMind (Cui et al., 2024a) paper, an agent-based check-in and security system can prompt a robot to approach and identify a person when the system cannot recognize them from low-resolution images captured by a ceiling-mounted camera or due to occlusion.
@@
Iterative Self-Refinement. For methods based on LLMs, due to their non-deterministic nature, incorporating contextual information into prompts does not always guarantee fully executable responses or complete satisfaction of personalization requirements. To address this, iterative self-refinement mechanism are proposed to transforms LLMs into a closed-loop system (Liu et al., 2024b; Cui et al., 2024a; Guo et al., 2024b; Xu et al., 2024b; Liu et al., 2024e; Yang et al., 2024b). Specifically, the LLM’s output (e.g., a plan generated by an LLM-agent) receives feedback from the environment or the user if the response is not executable, possibly due to overlooked environmental or user-specific factors. The LLM is then reprompted with information about the execution error or user feedback, generating a revised response. This process repeats until the generated response fully meets the environmental and personalization requirements.
@@ 6. Criterion III: Safety
Safety in IoT systems refers to preventing the system and its components from causing physical harm or posing threats, and to protecting the surrounding environment from such risks (Atlam and Wills, 2020). It is worth noting that in other related surveys (e.g., Ma et al. (2025)), safety may also encompass security, which refers to protecting models from external threats such as adversarial, jailbreak attacks. In this survey, we consider them differently as Shi et al. (2024a) and address security separately in Section 7.
@@
Figure 3. Illustrative example of five safety-enhancing methods introduced in this paper for applying LLMs to IoT tasks, demonstrated through an autonomous driving scenario.
All studies reviewed in this survey focus on enhancing the safety of LLMs; therefore, we summarize the commonly used techniques for improving LLM safety in the context of IoT integration. These methods include: (1) safety requirement incorporation, (2) user request verification, (3) post-response verification, (4) interpretable decision-making, and (5) iterative self-refinement. Table 5 provides an overview of these methods, including their descriptions and the paradigms (prompt-based, agent-based, training-based) to which they are mostly commonly applied. Figure 3 presents an example of an LLM-controlled autonomous driving system to illustrate these safety enhancement methods.
Table 5. Summary of Safety-Enhancing Methods, with Descriptions and Typically Applicable Paradigms (Prompt-Based, Agent-Based, Training-Based).
Typically Applicable Paradigm
Method Description Prompt Agent Training
Safety Requirement Incorporation Embedding safety requirements into LLM inputs.
User Request Verification
Reasoning over both the user request and the current
environmental context to assess whether the task is safe
and compliant with predefined safety constraints.
Post-Response Verfication
Verifying that the output adheres to both general and
task-specific safety constraints.
Interpertable Decision-Making
Generating not only the final decision,
but also the reasoning procedure explicitly.
Iterative Self-Refinement Regenerating the output until satisfying the safety requirements.
@@
Safety Requirement Incorporation. Embedding safety requirements into LLM inputs, typically via system prompts, is a key strategy for promoting safety awareness (Xu et al., 2024b; Ravichandran et al., 2025; Brunke et al., 2025). This approach requires developers and domain experts to proactively identify potential risks and define appropriate constraints and mitigation strategies prior to deployment. In agent-based systems, where external information can be retrieved dynamically from online sources or databases, environment- and task-specific safety requirements can also be automatically incorporated into the prompt to ensure context-aware and adaptive safety compliance.
@@
User Request Verification. When presented with a task from user requests, an LLM should reason over both the request and the current environmental context to assess whether the task is safe and compliant with predefined safety constraints. Context-aware reasoning is essential, as high-level instructions may not be safe to execute across all situations (Cui et al., 2024b; Xu et al., 2024b; Ravichandran et al., 2025). For example, the model should recognize and reject unsafe user requests, such as driving at 200 miles per hour (322 km/h) on a snowy road.
@@
Post-Response Verification. After the LLM generates the response, it is crucial to verify that the output adheres to both general and task-specific safety constraints, either manually or automatically. Manual verification, however, can be time-consuming, especially when domain expertise and complex reasoning are required. Therefore, automated verification methods are essential (Xu et al., 2024b; Yang et al., 2024d; Quartey et al., 2024). For example, Xu et al. (2024b) utilizes the Z3 Python API to check safety constraints formulated as first-order logic (FOL) against LLM-generated plans, which are expressed in formal languages such as Linear Temporal Logic (LTL) (Xu et al., 2024b; Yang et al., 2024b). Alternatively, responses can be validated through simulations to ensure safe and correct execution before deployment in real-world environments.
@@
Interpretable Decision-Making. When the LLM makes the decisions or operates the physical devices, especially in the high-stake scenarios (e.g., healthcare, autonomous driving), interpretability of the final decision is very important. This means being able to provide not only the final decision, but also the reasoning procedure explicitly (e.g., reasoning chain in Chain-of-Thought). This can help humans identify if the generated decision is correct and avoid catastrophes if the LLM intermediate reasoning steps violates the safety rules (Cui et al., 2024b; Zhang et al., 2024b; Nouri et al., 2024).
@@
Iterative Self-Refinement. If an LLM output violates safety constraints, the model should be reprompted and required to regenerate the output iteratively, guided by feedback from either human reviewers or automated verification systems, until the generated response satisfies all predefined safety requirements (Xu et al., 2024b; Yang et al., 2024b, d). Alternatively, the LLM can generate initial outputs that are verified for safety, after which the model is iteratively fine-tuned on pairs of user instructions and verified outputs to improve its inherent safety awareness (Yang et al., 2024d).